Professor who has a mass of 95kg . Start down a sl
Answer #1

1)

The normal force acting on the box

N = mg cos 40

= (95kg) * 9.8 * cos 40

= 713.19 N

F = mgsin 40 - Ff

where Ff is force of kinetic friction = \mu * N

= 0.25 * 713.19

Ff = 178.29 N

F = mgsin 40 - Ff

ma = mgsin 40 - Ff

(95kg) * a= 95 * 9.8 *sin 40 - 178.29

a = 4.42 m/s2

2)

Velocity component of plane along x and y axes

Vx = (340km/hr) cos 12 = 332.57

Vy = (340km/hr) sin 12 = 70.68

Velocity component of wind along x and y axes

Vx = (40km/hr) cos 34 = 33.16

Vy = (40km/hr) sin 34 = 22.36

Resultant velocity components along x and y axes

Vrx = 332.57-33.16 = 299.41

Vry = 70.68 - 22.36 = 48.32

Vr = sqrt (Vrx2 + Vry2)

= 303.28 km/hr

Direction = tan-1 (Vrx / Vry)

= tan-1 (299.41/48.32)

= 80.83

3)

Distance travelled = 0.5 * g * t2

Distance travelled is height = 32m

32 = 0.5 * 9.8 * t2

t= 2.55 sec

Velocity = g * t

= 9.8 * 2.55

= 21.765 m/s

4)

Initial velocity, Vi = 34.9m/s

Angle, \theta = 35

For this question, the vertical velocity must be found to find the total time

Viy = vi sinθ

Viy = 34.9 m/s sin35°

Viy = 20.02 m/s

t= Viy/g

= 20.02 / 9.8

=2.042 sec

b)Horizontal displacement:

Vix = Vi cosθ

= 34.9 m/s cos35°

Vix = 28.58 m/s

dx = Vxt

dx = (28.58m/s)(2.042s)

dx = 58.36m

c) Peak height

With viy, dy can be determined using the following formula (remember that vfy is 0 m/s at maximum height!!!) :

vfy2 = viy2 – 2gdy

(0m/s)2 = (20.02 m/s)2 – 2(9.81 m/s2)dy

2 * 9.81 dy = 20.022

dy = 20.42m